OpenStudy (anonymous):

Solve the quadratic equation. (x + 3)^2 = 64

OpenStudy (shamil98):

Take the square root of both sides what do you get?

OpenStudy (anonymous):

take the square root of both sides to get rid of the exponents and then solving for x should be easy

OpenStudy (shamil98):

\[\Huge \sqrt{(x+3)^2} = \sqrt{64} \] ?

OpenStudy (anonymous):

x + 3 = 8 => x = 11

OpenStudy (shamil98):

@ganibl that is wrong.

OpenStudy (anonymous):

Sorry it should be x = +/- 8

OpenStudy (shamil98):

\[\Huge x+3= \pm 8 \] \[\Huge x = 5 , x = -11\]